A 59.2-kg person, running horizontally with a velocity of +3.63 m/s, jumps onto a 12.4-kg sled that is initially at rest.
a) Ignoring the effects of friction during the collision, find the velocity of the sled and person as they move away.

b)The sled and person coast 31.5 m on level snow before coming to rest. How strong is the force of kinetic friction between the sled and the snow?
1
Expert's answer
2020-03-20T09:38:25-0400

a) From the conservation of energy:


"Mv=(m+M)u"

"(59.2)(3.63)=(59.2+12.4)u"

"u=3.00\frac{m}{s}"

b)


"u^2=2\mu gd"

"3^2=2\mu (9.8)(31.5)"

"\mu=0.0146"